OpenStudy (anonymous):

The minimum value of \[x^{2}+2xy+3y^{2}-6x-2y\]where\[x,y \epsilon R\] is ???????

OpenStudy (anonymous):

can we use calculus ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

we are not of that level

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well i want to use quadratic equations i think u are more comfortable with them

OpenStudy (anonymous):

first let \(x^2+2xy+3y^2-6x-2y=z \) we want to find the minimum of z

OpenStudy (anonymous):

i want to arrange the equation to appear as a quadratic equation in terms of x

OpenStudy (anonymous):

so i have \( x^2+(2y-6)x+3y^2-2y-z=0 \)

OpenStudy (anonymous):

in three variables?

OpenStudy (anonymous):

no problem with that

OpenStudy (anonymous):

x is real so the discriminant of that quadratic equation must be greater than or equal to zero

OpenStudy (anonymous):

hence \( (2y-6)^2-4(3y^2-2y-z) \ge 0 \) simplify it u will get \( 2y^2+4y-9-t \ge 0 \)

OpenStudy (anonymous):

sorry u will get \( 2y^2+4y-9-t \le 0 \)

OpenStudy (anonymous):

right till here?

OpenStudy (anonymous):

oh waht a mistake its z not t

OpenStudy (anonymous):

yes then

OpenStudy (anonymous):

now we must have D>=0 for f(y)=2y^2+4y-9-z it will gives z>=-11 so the minimum is -11

OpenStudy (anonymous):

how -11?

OpenStudy (anonymous):

see if D<0 for f(y)=2y^2+4y-9-z then all of f(y) lies on the above of x-axis and then f(y)>0 so we must have D>=0

OpenStudy (anonymous):

D is discriminant

OpenStudy (anonymous):

how have you solved to -11?

OpenStudy (anonymous):

discriminant of 2y^2+4y-9-z is 16+8(z+9) so 16+8(z+9)>=0 or z>=-11

OpenStudy (anonymous):

interesting question; i've seen like this before

OpenStudy (anonymous):

pls explain

OpenStudy (anonymous):

why we have taken D >=0, we have to take D<=0

OpenStudy (anonymous):

ok see if D<=0 there is no real values for y and one of the conditions of question is to be y from real numbers

OpenStudy (anonymous):

ok understood

OpenStudy (anonymous):

If we would have taken the eq. in terms of x then also the ans. would have come -11??

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

but i am getting z >=-28 when taken it in terms of x.

OpenStudy (anonymous):

let me check it

OpenStudy (anonymous):

in terms of y

OpenStudy (anonymous):

in terms of x you have solved

OpenStudy (anonymous):

i got z>=-11 when we put it in terms of y in the very first equation

OpenStudy (anonymous):

so its allright

OpenStudy (anonymous):

what eq. you got in terms of y?

OpenStudy (anonymous):

32+3z+1>=0

OpenStudy (anonymous):

i got 2x^2-16x-1-z<=0

OpenStudy (anonymous):

i got 2x^2-16x-1-3z<=0

OpenStudy (anonymous):

ok thanks